# MedMamba: Vision Mamba for Medical Image Classification

### 06 Mar 2024 v1

@@ Abstract
Medical image classification is a very fundamental and crucial task in the field of computer vision. These years, CNN-based and Transformer-based models are widely used in classifying various medical images. Unfortunately, The limitation of CNNs in long-range modeling capabilities prevent them from effectively extracting fine-grained features in medical images , while Transformers are hampered by their quadratic computational complexity. Recent research has shown that the state space model (SSM) represented by Mamba can efficiently model long-range interactions while maintaining linear computational complexity. Inspired by this, we propose Vision Mamba for medical image classification (MedMamba). More specifically, we introduce a novel Conv-SSM module, which combines the local feature extraction ability of convolutional layers with the ability of SSM to capture long-range dependency. To demonstrate the potential of MedMamba, we conduct extensive experiments using three publicly available medical datasets with different imaging techniques (i.e., Kvasir (endoscopic images), FETAL_PLANES_DB (ultrasound images) and Covid19-Pneumonia-Normal Chest X-Ray (X-ray images)) and two private datasets built by ourselves. Experimental results show that the proposed MedMamba performs well in detecting lesions in various medical images. To the best of our knowledge, this is the first Vision Mamba tailored for medical image classification. The purpose of this work is to establish a new baseline for medical image classification tasks and provide valuable insights for the future development of more efficient and effective SSM-based artificial intelligence algorithms and application systems in the medical. Source code has been available at https://github.com/YubiaoYue/MedMamba.
@@ 3.1 Datasets
Cervical lymph node lesion ultrasound images (Abbreviated as CLNNUS). CLNNUS is a private dataset containing 3392 cervical lymph node ultrasound images. Specifically, these images were obtained from 480 patients in the Ultrasound Department of the Second Affiliated Hospital of Guangzhou Medical University. The entire dataset is divided into four categories by clinical experts based on pathological biopsy results: normal lymph nodes (referred to as normal), benign lymph nodes (referred to as benign), malignant primary lymph nodes (referred to as primary), and malignant metastatic lymph nodes (referred to as metastatic). The number of normal, benign, primary and metastatic images are 1217, 601, 236 and 1338 respectively.
FETAL_PLANES_DB[27] (Abbreviated as FETAL). A large dataset of routinely acquired maternal-fetal screening ultrasound images collected from two different hospitals by several operators and ultrasound machines. All images were manually labeled by an expert maternal fetal clinician. Images are divided into 6 classes: four of the most widely used fetal anatomical planes (Abdomen, Brain, Femur and Thorax), the mother’s cervix (widely used for prematurity screening) and a general category to include any other less common image plane. Fetal brain images are further categorized into the 3 most common fetal brain planes (Trans-thalamic, Trans-cerebellum, Trans-ventricular) to judge fine grain categorization performance. Based on FETAL’s metadata, we categorize it into six categories. The number of images for each category is as follows: Fetal abdomen (711 images), Fetal brain (3092 images), Fetal femur (1040 images), Fetal thorax (1718 images), Maternal cervis (1626 images), and Other (4213 images) (dataset url: https://zenodo.org/records/3904280).
Covid19-Pneumonia-Normal Chest X-Ray Images[28, 29] (Abbreviated as CPN-CX). Shastri et al collected a large number of publicly available and domain recognized X-ray images from the Internet, resulting in CPN-CX. The CPN-CX dataset is divided into 3 categories, namely COVID, NORMAL and PNEUMONIA. All images are preprocessed and resized to 256x256 in PNG format. It helps the researcher and medical community to detect and classify COVID19 and Pneumonia from Chest X-Ray Images using Deep Learning (dataset url: https://data.mendeley.com/datasets/dvntn9yhd2/1).
@@ 3.4 Main results and findings
In previous work, researchers and developers tended to use classification models carefully designed for ImageNet[32] to classify medical images. Although these models were initially applied to natural image classification, experimental results have shown that they are also effective in medical image classification. Therefore, to demonstrate the potential of MedMamba, we compared it with some state-of-the-art models applied in medical image classification tasks, and the experimental results are shown in Table 1. According to the results, although the proposed MedMamba architecture is very straightforward, it still performs well in medical image classification tasks and surpasses CNNs-based and Transformers-based models in some datasets.

### 21 Mar 2024 v2

@@ Abstract
Medical image classification is a very fundamental and crucial task in the field of computer vision. These years, CNN-based and Transformer-based models have been widely used to classify various medical images. Unfortunately, The limitation of CNNs in long-range modeling capabilities prevents them from effectively extracting features in medical images, while Transformers are hampered by their quadratic computational complexity. Recent research has shown that the state space model (SSM) represented by Mamba can efficiently model long-range interactions while maintaining linear computational complexity. Inspired by this, we propose Vision Mamba for medical image classification (MedMamba). More specifically, we introduce a novel Conv-SSM module. Conv-SSM combines the local feature extraction ability of convolutional layers with the ability of SSM to capture long-range dependency, thereby modeling medical images with different modalities. To demonstrate the potential of MedMamba, we conducted extensive experiments using 14 publicly available medical datasets with different imaging techniques and two private datasets built by ourselves. Extensive experimental results demonstrate that the proposed MedMamba performs well in detecting lesions in various medical images. To the best of our knowledge, this is the first Vision Mamba tailored for medical image classification. The purpose of this work is to establish a new baseline for medical image classification tasks and provide valuable insights for the future development of more efficient and effective SSM-based artificial intelligence algorithms and application systems in the medical. Source code has been available at https://github.com/YubiaoYue/MedMamba.
@@ 3.1 Datasets
Figure 2: Sample images of medical images with different modality.
Cervical lymph node lesion ultrasound images (Abbreviated as Cervical-US). Cervical-US is a private dataset containing 3392 cervical lymph node ultrasound images. Specifically, these images were obtained from 480 patients in the Ultrasound Department of the Second Affiliated Hospital of Guangzhou Medical University. The entire dataset is divided into four categories by clinical experts based on pathological biopsy results: normal lymph nodes (referred to as normal), benign lymph nodes (referred to as benign), malignant primary lymph nodes (referred to as primary), and malignant metastatic lymph nodes (referred to as metastatic). The number of normal, benign, primary and metastatic images are 1217, 601, 236 and 1338 respectively.
FETAL_PLANES_DB[27] (Abbreviated as Fetal-US). A large dataset of routinely acquired maternal-fetal screening ultrasound images collected from two different hospitals by several operators and ultrasound machines. All images were manually labeled by an expert maternal fetal clinician. Images are divided into 6 classes: four of the most widely used fetal anatomical planes (Abdomen, Brain, Femur and Thorax), the mother’s cervix (widely used for prematurity screening) and a general category to include any other less common image plane. Fetal brain images are further categorized into the 3 most common fetal brain planes (Trans-thalamic, Trans-cerebellum, Trans-ventricular) to judge fine grain categorization performance. Based on FETAL’s metadata, we categorize it into six categories. The number of images for each category is as follows: Fetal abdomen (711 images), Fetal brain (3092 images), Fetal femur (1040 images), Fetal thorax (1718 images), Maternal cervis (1626 images), and Other (4213 images) (dataset url: https://zenodo.org/records/3904280).
Covid19-Pneumonia-Normal Chest X-Ray Images[28, 29] (Abbreviated as CPN-Xray). Shastri et al collected a large number of publicly available and domain recognized X-ray images from the Internet, resulting in CPN-Xray. The CPN-Xray dataset is divided into 3 categories, namely COVID, NORMAL and PNEUMONIA. All images are preprocessed and resized to 256x256 in PNG format. It helps the researcher and medical community to detect and classify COVID19 and Pneumonia from Chest X-Ray Images using Deep Learning (dataset url: https://data.mendeley.com/datasets/dvntn9yhd2/1).
@@
Medmnist2D. MedMNIST2D datasets include a set of 12 pre-processed datasets that include CT, X-ray, ultrasound and OCT images. These datasets are used in various classification tasks including multi-label, ordinal, multi-class, regression, and binary. The size of the data in this collection varies from at least 100 to more than 100,000. Due to the focus of this work on medical image classification, 10 datasets from Medmnist were selected. As shown in Table 1, the diversity of these datasets has created a favorable criterion for classification tasks. The pre-processing and split of the datasets into training, validation and test subsets have been done according to [32, 33] (dataset url: https://medmnist.com/).
Table 1: A brief overview of the 10 datasets selected from Medmnist2D. MC: Multi-Class. BC: Binary-Class.
Name Data Modality Task (# Classes / Labels) # Samples # Training / Validation / Test
MedMNIST2D
PathMNIST Colon Pathology MC (9) 107,180 89,996 / 10,004 / 7,180
OCTMNIST Retinal OCT MC (4) 109,309 97,477 / 10,832 / 1,000
DermaMNIST Dermatoscope MC (7) 10,015 7,007 / 1,003 / 2,005
PneumoniaMNIST Chest X-Ray BC (2) 5,856 4,708 / 524 / 624
BreastMNIST Breast Ultrasound BC (2) 780 546 / 78 / 156
TissueMNIST Kidney Cortex Microscope MC (8) 236,386 165,466 / 23,640 / 47,280
BloodMNIST Blood Cell Microscope MC (8) 17,092 11,959 / 1,712 / 3,421
OrganAMNIST Abdominal CT MC (11) 58,850 34,581 / 6,491 / 17,778
OrganCMNIST Abdominal CT MC (11) 23,660 13,000 / 2,392 / 8,268
OrganSMNIST Abdominal CT MC (11) 25,221 13,940 / 2,452 / 8,829
@@ 3.4 Main results and findings
In previous work, researchers and developers tended to use classification models carefully designed for ImageNet[34] to classify medical images. Although these models were initially applied to natural image classification, experimental results have shown that they are also effective in medical image classification. Therefore, to demonstrate the potential of MedMamba, we compared it with some state-of-the-art models applied in medical image classification tasks, and the experimental results are shown in Table 2 and Table3. According to the results, although the proposed MedMamba architecture is very straightforward, it still performs well in medical image classification tasks and surpasses CNNs-based and Transformers-based models in some datasets.
@@
Methods PathMNIST DermaMNIST OCTMNIST PneumoniaMNIST BreastMNIST
AUC ACC AUC ACC AUC ACC AUC ACC AUC ACC
ResNet18 (28)[35] 0.983 0.907 0.917 0.735 0.943 0.743 0.944 0.854 0.901 0.863
ResNet18 (224)[35] 0.989 0.909 0.920 0.754 0.958 0.763 0.956 0.864 0.891 0.833
ResNet50 (28)[35] 0.990 0.911 0.913 0.735 0.952 0.762 0.948 0.854 0.857 0.812
ResNet50 (224)[35] 0.989 0.892 0.912 0.731 0.958 0.776 0.962 0.884 0.866 0.842
auto-sklearn[40] 0.934 0.716 0.902 0.719 0.887 0.601 0.942 0.855 0.836 0.803
AutoKeras[41] 0.959 0.834 0.915 0.749 0.955 0.763 0.947 0.878 0.871 0.831
Google AutoML[42] 0.944 0.728 0.914 0.768 0.963 0.771 0.991 0.946 0.919 0.861
MedVit-T[43] 0.994 0.938 0.914 0.768 0.961 0.767 0.993 0.949 0.934 0.896
MedVit-S[43] 0.993 0.942 0.937 0.780 0.960 0.782 0.995 0.961 0.938 0.897
MedVit-L[43] 0.984 0.933 0.920 0.773 0.945 0.761 0.991 0.921 0.929 0.883
MedMamba 0.997 0.951 0.907 0.758 0.993 0.914 0.965 0.912 0.879 0.872
Methods BloddMNIST TissueMNIST OrganAMNIST OrganCMNIST OrganSMNIST
AUC ACC AUC ACC AUC ACC AUC ACC AUC ACC
ResNet18 (28)[35] 0.998 0.958 0.930 0.676 0.997 0.935 0.992 0.900 0.972 0.782
ResNet18 (224)[35] 0.998 0.963 0.933 0.681 0.998 0.951 0.994 0.920 0.974 0.778
ResNet50 (28)[35] 0.997 0.956 0.931 0.680 0.997 0.935 0.992 0.905 0.972 0.770
ResNet50 (224)[35] 0.997 0.950 0.932 0.680 0.998 0.947 0.993 0.911 0.975 0.785
auto-sklearn[40] 0.984 0.878 0.828 0.532 0.963 0.762 0.976 0.829 0.945 0.672
AutoKeras[41] 0.998 0.961 0.941 0.703 0.994 0.905 0.990 0.879 0.974 0.813
Google AutoML[42] 0.998 0.966 0.924 0.673 0.990 0.886 0.988 0.877 0.964 0.749
MedVit-T[43] 0.996 0.950 0.943 0.703 0.995 0.931 0.991 0.901 0.972 0.789
MedVit-S[43] 0.997 0.951 0.952 0.731 0.996 0.928 0.993 0.916 0.987 0.805
MedVit-L[43] 0.996 0.954 0.935 0.699 0.997 0.943 0.994 0.922 0.973 0.806
MedMamba 0.999 0.984 0.930 0.694 0.998 0.954 0.995 0.924 0.978 0.808